OpenStudy (anonymous):

suppose z varies directly as the square of x and inversely as y. if z=8 when x=4 and y=6, find z when x=6 and y=12

OpenStudy (saifoo.khan):

\[\Huge z = \frac{kx^2}{y}\]

OpenStudy (saifoo.khan):

find k in the first step!

OpenStudy (saifoo.khan):

\[\large 8 = \frac{4(k)}{6}\]

OpenStudy (saifoo.khan):

can u do it now?

OpenStudy (anonymous):

yea, thanks

OpenStudy (saifoo.khan):

welks!

OpenStudy (anonymous):

\[Z=C*(x ^{2}/y)\] 8=C(16/6) FIND C C=3 then Z=3∗(x2/y) Z=3∗(36/12) Z=9